Consider the following reaction occurring at 25 de
Answer #1

A.) Spontaneous reaction can be dtermined by calculating delta G ( \DeltaGo)

\DeltaGo= \sum\DeltaG products - \sum\DeltaGreactants = 2(-204.9 ) - ( -394.4) + (-62.3) = -409.8 - (-456.7) = 46.9 kJ/mol

The reaction is non spontaneous because delta G is postive

B.) ln Keq = - \DeltaGo/ RT = -46.9 kJ/mol / (0.008314kJmol^-1 K^-1 ) * ( 373K) = -15.12

ln Keq = -15.12

Keq = 2.713 x 10^-7

\DeltaHo = \sum\DeltaH products - \sum\DeltaHreactants = 2(-219.9) - (-95.7) + (-393.5) = - 439.8 -(-489.2) =49.4 kJ/mol

C.) \DeltaG = \DeltaH - T\DeltaS

\DeltaG= 0

T = \DeltaH / \DeltaS = 49.4/46.9 kJ/mol = 1.05 oC